ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 243, 2021

Adin, R. M., I. M. Gessel, V. Reiner and Y. Roichman Cyclic quasi-symmetric functions

437

Aguilera, J. P. Long Borel Games  273

Balakrishnan, J. S., A. Besser, F. Bianchi and J. S. Müller Explicit quadratic Chabauty over number fields

185

Basso, G. Almost minimal orthogonal projections  355
Bergweiler, W. and J. Ding Non-escaping points of Zorich maps  27

Besser, A. See J. S. Balakrishnan Bianchi, F. See J. S. Balakrishnan Bonomo, W., H. Lima and P. Varandas The rotation sets of most

volume preserving homeomorphisms on 𝕋^{d} are stable, convex and rational polyhedrons

81

Danilenko, A. I. On the bounded cohomology for ergodic nonsingular actions of amenable groups

421

Ding, J. See W. Bergweiler Dinur, I., S. Khot, G. Kindler, D. Minzer and M. Safra On

non-optimally expanding sets in Grassmann graphs

377

García-Ramos, F., T. Jäger and X. D. Ye Mean equicontinuity, almost automorphy and regularity

155

Gessel, I. M. See R. M. Adin Hibi, T., M. Lasoń, K. Matsuda, M. Michałek and M. Vodička

Gorenstein graphic matroids

1

Huisgen-Zimmermann, B. Dualities from iterated tilting  315

Jäger, T. See F. García-Ramos Keevash, P., J. Long, B. Narayanan and A. Scott A universal

exponent for homeomorphs

141

Khot, S. See I. Dinur Kindler, G. See I. Dinur

Kumashiro, S. Ideals of reduction number two  45

Lasoń, M. See T. Hibi Lima, H. See W. Bonomo Long, J. See P. Keevash Luo, J., X. Wang and Q.-S. Wu Poisson Dixmier–Moeglin equiva-

lence from a topological point of view 103 Matsuda, K. See T. Hibi Michałek, M. See T. Hibi Minzer, D. See I. Dinur Müller, J. S. See J. S. Balakrishnan Narayanan, B. See P. Keevash Reiner, V. See R. M. Adin Roichman, Y. See R. M. Adin Safra, M. See I. Dinur Sauermann, L. On the size of subsets of 𝔽_{p}^{n} without p distinct ele-

ments summing to zero

63

Scott, A. See P. Keevash

Sodomaco, L. On the product of the singular values of a binary tensor  233

Varandas, P. See W. Bonomo Vodička, M. See T. Hibi Wang, X. See J. Luo Wu, Q.-S. See J. Luo Ye, X. D. See F. García-Ramos